# The WINTER Observatory: A One-Degree InGaAs Survey Camera to study the Transient Infrared Sky

### v1

@@ Abstract
The Wide-field Infrared Transient Explorer (WINTER) is a near-infrared time-domain survey instrument installed on a dedicated 1-meter robotic telescope at Palomar Observatory. The project takes advantage of the recent technology advances in time-domain astronomy, robotic telescopes, large-format sensors, and rapid data reduction and alert software for timely follow up of events. Since June of 2023, WINTER robotically surveys the sky each night to a median depth of mag, balancing a variety of science programs including searching for kilonovae from gravitational-wave alerts, blind surveys to study galactic and extragalactic transients and variables, and building up reference images of the near-infrared sky. The project also serves as a technology demonstration for new large-format Indium Gallium Arsenide (InGaAs) sensors for wide-field science in the near infrared without cryogenically cooled optics or detectors. WINTER’s custom camera combines six InGaAs sensors with a novel tiled fly’s-eye optical design to cover a 1 degree-squared field of view with 1 arcsecond pixels in the -, -, and shortened--band filters (0.9 - 1.7m). This paper presents the design, performance, and early on-sky science of the WINTER observatory.
@@ I Introduction
Despite the recent rise in optical time-domain surveys, such as ATLAS (atlas), DeCAM (DECAM), Pan-STARRS (pan), ZTF (ztf2019), and GOTO (goto) and the upcoming Rubin Telescope’s LSST survey (Rubin), the near-infrared time-domain sky remains relatively unexplored. High sky backgrounds and the cost of the state-of-the-art HgCdTe (Mercury Cadmium Telluride) sensors prevent similar growth in near-infrared time-domain astronomy. Legacy and contemporary near-infrared surveys—for example, TMSS (tmss), 2MASS (2mass), UKIRT (UKIRT), VISTA (VISTA), WISE/NEOWISE (wise), and Palomar Gattini IR (De:2019xhw)—often trade off between coverage and depth, resulting in either relatively shallow surveys (for example, Palomar Gattini IR, mag; De:2019xhw) or are restricted to small areas on sky (such as the UKIRT Infrared Deep Sky Survey (UKIDSS) which covered 7,500 square degrees (UKIRT) or the VISTA Variables in the Via Lactea (VVV) survey covering 520 square degrees (VISTA)). Consequently, in near-infrared wavebands, there is a gap in deep, all-sky reference images and characterization of near-infrared transients which would benefit the time-domain community and lay groundwork for the upcoming Nancy Grace Roman Space Telescope.
@@
The Wide-Field Infrared Transient Explorer (WINTER) is a new near-infrared time-domain survey instrument on a dedicated 1-meter robotic telescope at Palomar Observatory (Lourie:2020; Frostig:2020). WINTER covers over 1 square degree with a 90% fill factor in the (1.0 m), (1.2 m) and shortened -bands (1.6 m). The instrument was commissioned in June of 2023 and is operating robotically each night, surveying around 400 degrees per week to an average depth of magnitudes.
@@
WINTER is designed around three primary goals: (1) rapid robotic follow-up of gravitational-wave alerts to search for kilonovae, (2) near-infrared time-domain surveys, and (3) general observing support for the WINTER community (Frostig:2020). The near-infrared colors uniquely study the r-process materials produced in kilonovae, allowing two weeks or more to search for electromagnetic signatures from the gravitational-wave events (Frostig:2022). Beyond kilonovae, near-infrared observations are well suited for dust-obscured, cool, and intrinsically red transients, variables, and static sky science. Between following up gravitational-wave alerts, WINTER scans the accessible northern sky in a series of Galactic and extragalactic surveys studying a range of science including transiting exoplanets around low-mass stars, tidal disruption events, supernovae, stellar mergers, and young star outbursts (See Section II for a detailed introduction to WINTER science cases). Over a ten-year nominal lifespan, WINTER will also build up a deep, co-added image of the near-infrared sky, preparing for future infrared surveys and studying galactic structure, high-redshift quasi-stellar objects, and brown dwarfs.
@@
In addition to WINTER’s science drivers, the instrument also demonstrates the feasibility of Indium Gallium Arsenide (InGaAs) hybridized focal plane arrays as a cheaper alternative to HgCdTe sensors for near-infrared astronomy. Unlike HgCdTe, InGaAs sensors do not require cryogenic cooling to achieve background limited performance from the ground, further reducing the cost and complexity of the project. A prototype camera demonstrated this approach with InGaAs sensors and achieved sky-background limited photometry while running around -40∘C with a thermoelectric cooler (TEC) and secondary liquid cooling (Simcoe:2019aps; Sullivan:2013). WINTER also features a novel “fly’s-eye optical” design to fill the 1 square degree field of view with six non-buttable sensors with only minimal gaps, and has custom readout electronics controlled by field programmable gate arrays (FPGAs), enabling fine control over the highly-programmable sensors.
WINTER is leading a new class of InGaAs-based astronomical surveys, which currently includes the upcoming DREAMS telescope in the Southern Hemisphere (DREAMS). WINTER, DREAMS, and the newly operating HgCdTe-based PRime-focus Infrared Microlensing Experiment (PRIME; PRIME_performance; PRIME_camera; PRIME_optics), are at the head of a new effort to expand our knowledge of the near-infrared time-domain sky.
@@
This paper describes the WINTER instrument and early performance since its installation in June 2023. A more detailed treatment on WINTER’s science cases can be found in Section II. Section III describes the InGaAs sensors and their performance, along with WINTER’s custom readout electronics, firmware, and software for controlling the sensors. The instrument design is described in Section IV, including the camera’s novel fly’s-eye optical design, the mechanical design, optomechanics, and cooling of the instrument. Section V describes the observatory and 1-meter robotic telescope and Section VI outlines the robotic control, scheduling, and data reduction software. Finally, the paper discusses on-sky performance including a brief description of current surveys and preliminary results from select early science investigations (Section VII), and a conclusion (Section VIII).
@@ II Science Objectives
WINTER’s design directly flows down from its science goals: including the need for rapid robotic follow up with low overheads, near-infrared coverage, and a wide field of view (Frostig:2020). The science cases highlighted in this section introduce the motivation behind building the instrument, but are not a comprehensive list.
@@
Near-infrared observations are uniquely suited for studying kilonovae, thermal transients powered by rapid neutron-capture (r-process) nucleosynthesis in the neutron-rich ejecta from binary neutron star (BNS) or neutron star black hole (NSBH) mergers. The landmark BNS event, GW170817, detected in both gravitational waves and across the electromagnetic spectrum, demonstrated how heavy element synthesis in ejecta abundant in r-process materials blankets optical wavelengths, shifting the afterglow into the NIR (TheLIGOScientific:2017qsa; GBM:2017lvd; Kasen:2017sxr). Unlike optical emissions, which are model-predicted to be short-lived (under one week) and highly viewing-angle dependent, NIR emission is both long-lived (over one week) and isotropic, making it a promising kilonova signature for BNS and some NSBH mergers (Kasen:2013xka; Barnes:2016umi). Moreover, NIR surveys could detect up to 8–10 times more kilonovae than optical surveys (Zhu:2020ffa).
WINTER’s design directly supports kilonova discovery through several key features: (1) sensitivity to the Lanthanide-rich kilonova emission peak in the near-infrared; (2) a wide field of view; (3) fast robotic follow-up with low overhead resulting in the ability to quickly tile gravitational-wave uncertainty contours. The need for efficient BNS merger follow-up, which yielded no detections during the third gravitational-wave observing run (O3), informed WINTER’s project requirements, including the optical design and ensquared energy criteria, the noise budget and instrument cooling, and optical performance which minimize false positives in wide-field searches (Frostig:2020).
Frostig:2022 outlines an end-to-end simulation of WINTER follow-up observing of BNS mergers during the O4. The study predicts NIR kilonovae are longer lived and red kilonovae are detected 1.5 times further in the infrared than in the optical. The study provides a detailed simulation of WINTER’s projected ability to discover new kilonovae, estimating up to five new discoveries in O4. However, the sensitivities of both WINTER and the international network of gravitational wave detectors were lower than expected throughout O4: no BNS mergers were detected by LIGO, and despite several international follow-up campaigns of BBH and BH-NS mergers, there were no confirmed optical counterparts of GW events (e.g., Pillas:2025, including WINTER data).
@@ II.2 Peering through the dust
Heavily obscured regions can conceal infrared transients and variable sources. Programs like the SPitzer InfraRed Intensive Transients Survey (SPIRITs), which surveyed 200 nearby galaxies, identified 2457 variables and 78 transients over four years, with 60 transients too red for optical detection, suggesting that optical surveys may miss up to 40% of core-collapse supernovae (CCSNe) in dusty environments (Kasliwal_SPIRITs_2018; Jencson_2019). SPIRITs also uncovered a new class of intermediate-luminosity transients, bridging the gap between novae and supernovae (Kasliwal_SPIRITs_2017). Infrared time-domain observations also reveal variability in young stellar objects (YSOs) across diverse environments; for instance, Spitzer found that 50% of YSOs in regions like the Orion Nebula and Large Magellanic Cloud are variable, enabling studies of jets, accretion disks, and circumstellar matter (yso_mag; yso_orion; yso_var). Infrared surveys are equally important for mapping obscured Galactic structures, as demonstrated by the VVV survey, which uncovered previously unknown variable stars and star-forming regions, while UKIDSS expanded our knowledge of faint stellar populations in the Galactic plane (VISTA; UKIRT).
@@ II.3 Studying cool objects
Infrared observations are particularly effective for studying cool objects such as brown dwarfs, planets, red giants, and asymptotic giant branch (AGB) stars, as these objects have lower temperatures (below 300 K for brown dwarfs up to 5,000 K for red giants) that peak in infrared wavelengths (Faherty; Marocco). Infrared time-domain surveys have significantly advanced our understanding of red giants and AGB stars, revealing pulsation periods and episodic mass-loss events observable through dust production and circumstellar changes (Liljegren). Long-Period Variables (LPVs), including Miras and semi-regular AGB stars, display complex variability patterns, which WISE and VVV surveys have explored in depth (Groenewegen; Nikzat). For brown dwarfs, Spitzer’s time-resolved spectroscopy and photometry have shown rotational modulations related to cloud patterns and temperature variations, enriching atmospheric models (Metchev; Artigau). Infrared studies also illuminate cool object binaries, where interactions with compact companions, such as in symbiotic systems, produce infrared variability through accretion processes, jet formation, and dust in outflows (Miko). Furthermore, exoplanet studies around M-dwarfs, the most common stars in the Milky Way, benefit from infrared observations, as these cool, low-mass stars offer favorable conditions for detecting transits in red optical to infrared bands (Dressing; Charbonneau). For example, the WINTER prototype InGaAs camera has observed exoplanet transits at Las Campanas Observatory, showcasing the potential of infrared tools for exoplanet discovery (Simcoe:2019aps).
@@ II.4 Static sky science
Throughout the course of WINTER’s nominal ten-year lifespan, the project aims to build up a deep reference image of the near-infrared northern sky. In addition to facilitating transient detection, utilizing this reference image to find new sources in nightly WINTER images, the reference image will help build a near-infrared map of L and T dwarfs, quasars, and galactic structure while also providing a reference image for the broader community. The static reference image and a long-baseline time-domain survey will also provide valuable preparation for the upcoming time-domain surveys from the Vera C. Rubin Observatory and Nancy Grace Roman Space Telescope.
@@ III.1 InGaAs sensors
The six InGaAs sensors central to WINTER are custom AP1020 focal plane arrays (FPAs) developed in collaboration with Teledyne-FLIR and Sensor Creations (now Attollo Engineering Inc./Safran Defense & Space Inc.) for our program (Figure 2). With a high-definition format of 1920 by 1080 pixels and a 15m pitch, the sensors are the largest format InGaAs sensors ever manufactured for commercial use. InGaAs exhibits dark current 2–3 orders of magnitude lower than 1.7 m cutoff HgCdTe at equivalent temperature (Seshadri_2007_MCT_vs_ingaas_JPL_2007) down to C, or in the 170–300 K range. For ground-based imaging applications like WINTER, adequate dark current to reach background limited performance can thus be reached without cryogenic cooling, and the 1.7 m cutoff reduces sensitivity to thermal backgrounds compared to detectors like 2.5 m cutoff HgCdTe which require cryogenically cooled optics.
The hybridized complementary metal oxide semiconductor (CMOS) focal plane arrays consist of photosensitive lattice-matched InGaAs PIN (p-type/insulator/n-type) diodes hybridized to a readout integrated circuit (ROIC) with a capacitive transimpedance amplifier (CTIA) pixel architecture. With a lower-energy bandgap than silicon, InGaAs is sensitive in the near/shortwave-infrared (NIR/SWIR) between 0.9 m and its 1.7 m room-temperature direct bandgap cutoff wavelength. The highly-programmable ROIC allows for multiple imaging modes, field-programmable gain, and application-specific sensor tuning. In contrast to the source follower per detector (SF) readout common in astronomy detectors like the Teledyne Hawaii family, the CTIA architecture supports video rate imaging at the price of increased read noise (ROIC_pixel_architectures_guellec_2014). An optimized CTIA readout can also deliver a higher degree of linearity than SF. To help mitigate this increased read noise (see Section III.4.4), the AP1020 ROIC features options for on-chip correlated double sampling and multiple non-destructive reads which substantially reduce the effective read noise. The sensor read out with eight interleaved data output channels, allowing for full frame, continuous sampling at a frame rate of up to 30 Hz. The focal plane array is bonded to a two-stage thermoelectric cooler (TEC) in hermetically-sealed vacuum housing (see further details in Section III.3).

### v2

@@ Abstract
The Wide-field Infrared Transient Explorer (WINTER) is a near-infrared time-domain survey instrument operating on a dedicated 1-meter robotic telescope at Palomar Observatory. The project takes advantage of recent technology advances in time-domain astronomy, robotic telescopes, large-format sensors, and rapid data reduction and alert software for timely follow up of events. Since June of 2023, WINTER robotically surveys the sky each night to a median depth of mag, balancing a variety of science programs including searching for kilonovae from gravitational-wave alerts, blind surveys to study galactic and extragalactic transients and variables, and building up reference images of the near-infrared sky. The project also serves as a technology demonstration for new large-format Indium Gallium Arsenide (InGaAs) sensors for wide-field science in the near infrared without cryogenically cooled optics or detectors. WINTER’s custom camera combines six InGaAs sensors with a novel tiled fly’s-eye optical design to cover a 1 deg2 field of view with 1 arcsecond pixels in the -, -, and shortened--band filters (0.9 - 1.7m). This paper presents the design, performance, and early on-sky science of the WINTER observatory.
@@ I Introduction
Despite the recent rise in optical time-domain surveys, such as ATLAS (Tonry et al., 2018), DeCAM (Flaugher et al., 2015), Pan-STARRS (Kaiser et al., 2010), ZTF (Bellm et al., 2019b), and GOTO (Dyer et al., 2022) and the upcoming Rubin Telescope’s LSST survey (Ivezić et al., 2019), the near-infrared (near-IR) time-domain sky remains relatively unexplored. High sky backgrounds and the cost of the state-of-the-art HgCdTe (Mercury Cadmium Telluride) sensors prevent similar growth in near-IR time-domain astronomy. Legacy and contemporary near-IR surveys—for example, TMSS (Neugebauer and Leighton, 1969), 2MASS (Skrutskie et al., 2006), UKIRT (Lawrence et al., 2007), VISTA (Dalton et al., 2006), WISE/NEOWISE (Wright et al., 2010), and Palomar Gattini IR (De et al., 2020)—often trade off between coverage and depth, resulting in either relatively shallow surveys (for example, Palomar Gattini IR, mag; De et al. (2020)) or are restricted to small areas on sky (such as the UKIRT Infrared Deep Sky Survey (UKIDSS) which covered 7,500 deg2 (Lawrence et al., 2007) or the VISTA Variables in the Via Lactea (VVV) survey covering 520 deg2 (Dalton et al., 2006)). Consequently, in near-IR wavebands, there is a gap in deep, all-sky reference images and characterization of near-IR transients which would benefit the time-domain community and lay groundwork for the upcoming Nancy Grace Roman Space Telescope.
@@
The Wide-Field Infrared Transient Explorer (WINTER) is a new near-IR time-domain survey instrument on a dedicated 1-meter robotic telescope at Palomar Observatory (Lourie et al., 2020; Frostig et al., 2020). WINTER covers over 1 deg2 with a 90% fill factor in the (1.0 m), (1.2 m) and shortened -bands (1.6 m). The instrument was commissioned in June of 2023 and is operating robotically each night, surveying around 400 degrees per week to an average depth of magnitudes.
@@
WINTER is designed around three primary goals: (1) rapid robotic follow-up of gravitational-wave alerts to search for kilonovae, (2) near-IR time-domain surveys, and (3) general observing support for the WINTER community (Frostig et al., 2020). The near-IR colors uniquely study the r-process materials produced in kilonovae, allowing two weeks or more to search for electromagnetic signatures from gravitational-wave events (Frostig et al., 2022a). Beyond kilonovae, near-IR observations are well suited for dust-obscured, cool, and intrinsically red transients, variables, and static sky science. Between following up gravitational-wave alerts, WINTER scans the accessible northern sky in a series of Galactic and extragalactic surveys studying a range of science including transiting exoplanets around low-mass stars, tidal disruption events, supernovae, stellar mergers, and young star outbursts (See Section II for a detailed introduction to WINTER science cases). Over a ten-year nominal lifespan, WINTER will also build up a deep, co-added image of the near-IR sky, preparing for future infrared surveys and studying galactic structure, high-redshift quasi-stellar objects, and brown dwarfs.
@@
In addition to WINTER’s science drivers, the instrument demonstrates the feasibility of Indium Gallium Arsenide (InGaAs) hybridized focal plane arrays as a cheaper alternative to HgCdTe sensors for near-IR astronomy. Unlike HgCdTe, InGaAs sensors do not require cryogenic cooling to achieve background-limited performance from the ground, further reducing the cost and complexity of the project. A prototype camera demonstrated this approach with InGaAs sensors and achieved sky-background limited photometry while running around -40∘C with a thermoelectric cooler (TEC) and secondary liquid cooling (Simcoe et al., 2019; Sullivan et al., 2013). WINTER also features a novel “fly’s-eye optical” design to fill the 1 deg2 field of view with six non-buttable sensors with only minimal gaps, and has custom readout electronics controlled by field programmable gate arrays (FPGAs), enabling fine control over the highly-programmable sensors.
WINTER is part of a new class of InGaAs-based astronomical surveys. Parallel efforts in this category include the upcoming DREAMS telescope in the Southern Hemisphere (Travouillon et al., 2020) and the smaller Antarctic Infrared Binocular Telescope (AIRBT; Dong et al. 2024, 2025). Together with the newly operating HgCdTe-based PRime-focus Infrared Microlensing Experiment (PRIME; Durbak et al. 2024; Kutyrev et al. 2023; Yama et al. 2023), these projects reflect a broader effort to expand coverage of the near-IR time-domain sky. These instruments further prepare for upcoming large-scale surveys such as the Nancy Grace Roman Space Telescope (Spergel et al., 2015) and Cryoscope (Kasliwal et al., 2025).
@@
This paper describes the WINTER instrument and early performance since its installation in June 2023. A more detailed treatment on WINTER’s science cases can be found in Section II. Section III describes the InGaAs sensors and their performance, along with WINTER’s custom readout electronics, firmware, and software. The instrument design is described in Section IV, including the camera’s novel fly’s-eye optical design, the mechanical design, optomechanics, and cooling of the instrument. Section V describes the observatory and 1-m robotic telescope, and Section VI outlines the robotic control, scheduling, and data reduction software. Finally, the paper discusses on-sky performance, including a brief description of current surveys and preliminary results from select early science investigations (Section VII) and concludes (Section VIII).
@@ II Science Objectives
WINTER’s design flows directly down from its science goals: rapid robotic follow up with low overheads, near-IR coverage, and a wide field of view (Frostig et al., 2020). The science cases highlighted in this section introduce the motivation behind building the instrument, but are not a comprehensive list.
@@
Near-IR observations are uniquely suited for studying kilonovae, thermal transients powered by rapid neutron-capture (r-process) nucleosynthesis in the neutron-rich ejecta from binary neutron star (BNS) or neutron star black hole (NSBH) mergers. The landmark BNS event, GW170817, detected in both gravitational waves and across the electromagnetic spectrum, demonstrated that r-process-rich ejecta obscure optical wavelengths, shifting the afterglow into the near-IR (Abbott et al., 2017a, b; Kasen et al., 2017). Unlike optical emissions, which are predicted to be short-lived and highly viewing-angle dependent, near-IR emission is both long-lived (over one week) and isotropic, making it a promising kilonova signature for BNS and some NSBH mergers (Kasen et al., 2013; Barnes et al., 2016). Studies show near-IR surveys could detect up to 8–10 times more kilonovae than optical surveys (Zhu et al., 2021).
To enable kilonova discovery, WINTER combines several key features: (1) sensitivity to the Lanthanide-rich kilonova emission peak in the near-IR; (2) a wide field of view; (3) fast robotic follow-up with low overhead resulting in the ability to quickly tile gravitational-wave uncertainty contours.
Frostig et al. (2022a) presents an end-to-end simulation of the WINTER follow-up observing of BNS mergers during O4. The study predicts that near-IR kilonovae are longer lived and red kilonovae are detected 1.5 times farther in the IR than in the optical. Detailed simulations of WINTER observing predicted up to five kilonova discoveries during O4. However, the sensitivities of both WINTER and the international network of gravitational wave detectors were lower than expected throughout O4. No BNS mergers were detected by LIGO, and despite several international follow-up campaigns of BBH and BH-NS mergers, there were no confirmed optical counterparts of GW events (e.g., Pillas et al. 2025, including WINTER data).
@@ II.2 Peering through the dust
Heavily obscured regions can conceal IR transients and variable sources. For example, the SPitzer InfraRed Intensive Transients Survey (SPIRITs) surveyed 200 nearby galaxies and identified 2457 variables and 78 transients over four years. Sixty of those transients were too red for optical detection, suggesting that optical surveys may miss up to 40% of core-collapse supernovae in dusty environments (Kasliwal et al., 2018; Jencson et al., 2019). SPIRITs also uncovered a new class of intermediate-luminosity transients, bridging the gap between novae and supernovae (Kasliwal et al., 2017). IR monitoring also probes variability in young stellar objects (YSOs) across diverse environments. In the Orion Nebula and the Large Magellanic Cloud, Spitzer observations showed that approximately 50% of YSOs exhibit variability, enabling studies of jet activity, accretion processes, and circumstellar dust evolution (Megeath et al., 2012; Gruendl and Chu, 2009; Lee et al., 2024). IR surveys further trace dust-obscured Galactic structure. The VVV survey revealed previously unknown variable stars and star-forming regions in the Galactic plane, while UKIDSS extended sensitivity to fainter stellar populations in crowded and highly extincted fields (Dalton et al., 2006; Lawrence et al., 2007).
@@ II.3 Studying cool objects
IR observations are essential for characterizing cool objects such as brown dwarfs, planets, red giants, and asymptotic giant branch (AGB) stars. These objects emit most of their flux at IR wavelengths due to their lower effective temperatures—below 300 K for brown dwarfs and up to 5000 K for red giants (Faherty et al., 2016; Marocco et al., 2019). Time-domain IR surveys reveal pulsation periods and episodic mass loss in red giants and AGB stars, tracing dust production and circumstellar evolution (Liljegren et al., 2017). Long-period variables (LPVs), including Miras and semi-regular AGB stars, show complex light curves that surveys such as WISE and VVV have systematically characterized (Groenewegen, 2022; Nikzat et al., 2022).Spitzer monitored brown dwarfs with time-resolved spectroscopy and photometry, detecting rotational modulations caused by cloud structure and temperature variations. These observations inform models of substellar atmospheres (Metchev et al., 2015; Artigau et al., 2009). IR monitoring also tracks variability in cool-object binaries. Systems such as symbiotic stars produce IR variability through accretion, jet activity, and dusty outflows driven by interactions with compact companions (Mikołajewska, 2012).
Exoplanet transit surveys targeting M-dwarfs also benefit from IR observations. These cool, low-mass stars dominate the stellar population of the Milky Way and provide favorable conditions for detecting transits in red-optical and IR bands (Dressing and Charbonneau, 2015; Charbonneau et al., 2007). The WINTER prototype InGaAs camera demonstrated this capability by observing exoplanet transits at Las Campanas Observatory (Simcoe et al., 2019).
@@ II.4 Static sky science
Over its ten-year lifespan, WINTER will build a deep reference image of the near-IR northern sky. The reference image will aid transient detection, support nightly source identification, and contribute to mapping L and T dwarfs, quasars, and Galactic structure. The static reference image and a long-baseline time-domain survey will provide valuable preparation for the upcoming time-domain surveys from the Vera C. Rubin Observatory and Nancy Grace Roman Space Telescope.
@@ III.1 InGaAs sensors
The six InGaAs sensors central to WINTER are custom AP1020 focal plane arrays (FPAs) developed in collaboration with Teledyne-FLIR and Sensor Creations (now Attollo Engineering Inc./Safran Defense & Space Inc.). With a high-definition format of 1920 by 1080 pixels and a 15m pitch, the sensors are the largest format InGaAs sensors ever manufactured for commercial use (Figure 2). InGaAs exhibits dark current 2–3 orders of magnitude lower than 1.7 m cutoff HgCdTe at equivalent temperature (Seshadri et al., 2007) down to C, or in the 170–300 K range. For ground-based imaging applications like WINTER, InGaAs sensors achieve background-limited performance without the need for cryogenic cooling. Their 1.7 m cutoff also reduces sensitivity to thermal background emission, unlike 2.5 m HgCdTe detectors, which require cryogenically cooled optics.
The hybridized complementary metal oxide semiconductor (CMOS) focal plane arrays consist of photosensitive lattice-matched InGaAs PIN (p-type/insulator/n-type) diodes hybridized to a readout integrated circuit (ROIC) with a capacitive transimpedance amplifier (CTIA) pixel architecture. With a lower-energy bandgap than silicon, InGaAs is sensitive in the near/shortwave-infrared (NIR/SWIR) between 0.9 m and its 1.7 m room-temperature direct bandgap cutoff wavelength. The highly-programmable ROIC allows for multiple imaging modes, field-programmable gain, and application-specific sensor tuning. In contrast to the source follower per detector (SF) readout common in astronomy detectors like the Teledyne Hawaii family, the CTIA architecture supports video rate imaging at the price of increased read noise (Guellec et al., 2017). An optimized CTIA readout can also deliver a higher degree of linearity than SF. To help mitigate this increased read noise (see Section III.4.4), the AP1020 ROIC features options for on-chip correlated double sampling and multiple non-destructive reads which substantially reduce the effective read noise. The sensor read out with eight interleaved data output channels, allowing for full frame, continuous sampling at a frame rate of up to 30 Hz. The focal plane array is bonded to a two-stage thermoelectric cooler (TEC) in hermetically-sealed vacuum housing (see further details in Section III.3).